# Integrating Protein Sequence and Expression Level to Analysis Molecular Characterization of Breast Cancer Subtypes

### v1

@@ Abstract
Breast cancer’s complexity and variability pose significant challenges in understanding its progression and guiding effective treatment. This study aims to integrate protein sequence data with expression levels to improve the molecular characterization of breast cancer subtypes and predict clinical outcomes. Using ProtGPT2, a language model designed for protein sequences, we generated embeddings that capture the functional and structural properties of proteins sequence. These embeddings were integrated with protein expression level to form enriched biological representations, which were analyzed using machine learning methods like ensemble K-means for clustering and XGBoost for classification. Our approach enabled successful clustering of patients into biologically distinct groups and accurately predicted clinical outcomes such as survival and biomarkers status, achieving high performance metrics, notably an F1 score of 0.88 for survival and 0.87 for biomarkers status prediction. Analysis of feature importance highlighted key proteins like KMT2C, GCN1, and CLASP2, linked to hormone receptor and Human Epidermal Growth Factor Receptor 2 (HER2) expression, which play a role in tumor progression and patient outcomes, respectively. Furthermore, protein-protein interaction networks and correlation analyses revealed the interdependence of proteins that may influence breast cancer subtype behaviors. These findings suggest that integrating protein sequence and expression data provides valuable insights into tumor biology and has significant potential to enhance personalized treatment strategies in breast cancer care.
@@ 1 Introduction
Breast cancer is still a major global health issue, making up a significant number of cancer cases and deaths among women. Despite advancements in early detection and treatment, the complexity and diversity of breast cancer present challenges in understanding its progression, predicting outcomes, and creating effective treatment plans. The classification of breast cancer into various subtypes based on molecular and clinical features is a crucial aspect of guiding treatment decisions and improving patient outcomes. Mainly, the presence or absence of specific biological markers, such as Estrogen Receptor (ER), Progesterone Receptor (PR), and Human Epidermal Growth Factor Receptor 2 (HER2), serves as a foundation for categorizing breast cancer into luminal, HER2-enriched, or basal-like subtypes [1]. These markers are essential in determining how a tumor may respond to hormone therapy or other targeted treatments, making their accurate classification vital for personalized medicine [2].
@@
Recent advances in proteomics and genomics have significantly enhanced the molecular understanding of breast cancer. Proteomics, the large-scale study of proteins, adds a critical layer to cancer research since proteins are the functional entities in cells that carry out biological processes. In breast cancer, shifts in protein expression patterns can provide valuable insights into tumor behavior, disease progression, and potential therapeutic targets. For instance, ER, PR, or HER2 changes can indicate how aggressive a tumor might be and what treatment strategies may be most effective. Understanding these markers not only aids in diagnosis but also plays a crucial role in therapeutic decision-making [2].
@@
One of the key innovations in this work involves integrating protein expression data with the underlying protein sequences. Protein expression levels indicate the abundance of specific proteins within cells, which can directly influence cellular function and tumor behavior [3]. On the other hand, protein sequences provide insights into the proteins’ structure and potential functional domains, which could clear their role in biological processes, including cancer progression [4]. Combining these two layers of information, protein abundance, and sequence features, we can capture both quantitative and qualitative aspects of the proteins, leading to a more comprehensive understanding of their roles in breast cancer biology.
@@
We employed ProtGPT2 [5], an advanced large language model (LLM) based on the GPT-2 architecture, to facilitate the integration of protein sequence and expression data. ProtGPT2 is trained explicitly for protein sequences, enabling it to generate meaningful embeddings that represent proteins’ functional and structural properties. These embeddings provide a high-dimensional representation that captures complex details of each protein sequence, allowing the model to leverage both local and global features of the sequences [5]. Integrating these sequence embeddings with protein expression levels in the context of breast cancer allows us to create a biologically meaningful representation of each protein. This enriched representation can be used for various downstream analyses, including clustering patients into distinct molecular subgroups and classifying clinical outcomes such as survival status and biomarker expression [6]. Our approach, which incorporates machine learning methods like ensemble K-means for clustering and classification algorithms such as XGBoost and Random Forest, aims to uncover relationships between protein features and clinical characteristics, offering new insights into the grouping of breast cancer patients and potential predictive biomarkers. Furthermore, the importance of trust and explainability in using artificial intelligence (AI) for healthcare has been highlighted in recent studies, including [7] that explore mechanisms to bridge the gap between advanced AI technologies and clinical practice, especially in the breast cancer.
Multiple clustering approaches are available, each suitable for different types of data and analytical needs. A spectral community detection approach based on an energy function has been proposed as an effective way to find communities within complex datasets by leveraging spectral properties to minimize or optimize the energy criteria [8, 9, 10, 11]. This method is beneficial for data with non-convex boundaries. Classical clustering methods like K-means are famous for their simplicity and effectiveness. However, they may struggle with complex data structures, whereas hierarchical clustering offers insight into relationships at multiple levels of granularity through dendrograms. Ensemble clustering combines different clustering results to produce a more robust and stable output, making it advantageous for datasets with inherent noise or variability. For projects involving complex biological datasets, such as protein interactions or gene expression, a combination of spectral, hierarchical, K-means, and ensemble clustering can be highly effective, offering insights into tightly bound communities, hierarchical relationships, and consensus groupings that reflect various facets of the data. On the other hand, protein interactions in breast cancer draw parallels to interconnected networks in epidemic modeling [12], which helps conceptualize the interactions between proteins and pathways. Managing data inconsistencies and uncertainty on the protein expression level is challenging [13]. The studies [14, 15] proposed robust trajectory tracking methods and sliding mode control, which could benefit the inconsistencies and uncertainty of measuring the protein expression level.
@@ 2.1 Data Source
In our work, we utilized the dataset of 12,553 proteins from the TCGA proteomics and then extracted protein sequence data from the National Center for Biotechnology Information (NCBI) database [18]. The protein sequence data were retrieved using API calls based on the RefSeq IDs provided in the TCGA dataset. The extracted protein sequences were obtained in FASTA format, ensuring compatibility with further bioinformatic and proteomic analyses. This additional data collection allowed for integrating protein sequence information with the proteomic expression levels, enriching the dataset used in our study.
@@ 2.5 Classification
Different classification algorithms such as XGBoost and Random Forest are applied to classify the patient data to predict different aspects like survival of patients. One of these predictions is to predict the status of three biological markers, ER, PR status, and HER2 Final Status, which are essential in breast cancer classification. These markers help determine the type of breast cancer, guide treatment decisions, and predict patient outcomes. Each marker is a target variable, a multi-output classification problem. The $X_{m}$ is used to predict them as a multi-label. The Hamming loss metric measures the fraction of incorrectly predicted labels, providing an overall view of the model’s ability to classify each target accurately. Lower Hamming loss indicates better classification performance. The formula for Hamming loss is given by [22]:
@@ 3 Results
To evaluate the effectiveness and robustness of integrating protein sequence data with protein expression levels, we applied XGBoost to predict the survival outcomes of breast cancer patients. The survival status, defined as either "alive" or "deceased", was predicted based on a comprehensive feature set comprising the integrated protein sequence and expression level and relevant clinical data such as tumor size and patient age. The classification model demonstrated a strong performance, achieving an F1 score of 0.88. This result underscores the model’s ability to distinguish between patients who are likely to survive and those at a higher risk.
We employed multiple classification algorithms to evaluate the predictive capability of our integrated protein sequence and expression level in determining the status of breast cancer biomarkers, including ER, PR, and HER2 status. We treated the prediction of these three markers as a multi-output classification problem, where each marker represents an individual target to be predicted simultaneously. Table 1 compares various classification algorithms, showcasing the precision, recall, F1 score, and Hamming loss for each model. These metrics are reported as weighted averages, reflecting the model’s ability to predict multiple target labels across all patient samples effectively.
@@
Table 1: Precision, recall, and F1 score for predicting all markers as a multi-output classification problem based on different algorithms reported in the weighted average with a sample size of 24
The results in Table 1 show that XGBoost achieved the best overall performance, with an F1 score of 0.87 and balanced precision and recall values of 0.88 each. This indicates that XGBoost was highly influential in predicting the multi-output targets accurately, achieving strong generalization across all three biomarkers while maintaining a relatively low Hamming loss of 0.18. CatBoost and AdaBoost also performed well, with CatBoost showing a slight edge in precision and Hamming loss compared to XGBoost, suggesting that these gradient-boosting algorithms are particularly suited for complex, multi-target classification tasks involving biological data. Notably, models like K-Nearest Neighbors and Support Vector Machines displayed relatively lower performance, particularly in precision and F1 score. This difference might be due to the high-dimensional nature of the integrated data, which XGBoost and CatBoost were better able to manage, likely due to their sophisticated handling of complex, non-linear relationships within the dataset.
To further investigate which proteins were most influential in predicting the biomarker statuses, we analyzed the feature importance scores from the XGBoost model. The integration of protein expression levels and sequence information was used for each protein, allowing the model to leverage both the abundance of the proteins and their functional sequence-derived attributes. Table 2 lists the top 10 proteins, including critical proteins with the highest importance scores during classification.
@@
Table 2: Importance score values for top 10 important proteins
Among the proteins, KMT2C and GCN1 were the most influential proteins, with importance scores of 0.119 and 0.114, respectively. KMT2C is a known epigenetic regulator, and its involvement in breast cancer suggests a potential link to tumor behavior and biomarker status. Similarly, GCN1 and CLASP2 are proteins associated with cellular growth and response to environmental stress, indicating their potential roles in modulating breast cancer subtype behavior.
@@
Figure 1: Correlation matrix of the top 10 important proteins in the biomarkers prediction
Figure 2 presents the protein-protein interaction (PPI) network for the top 10 most important proteins identified from the XGBoost feature importance analysis. This network was constructed using STRING-db [23], a well-known database for retrieving known and predicted protein-protein interactions. The nodes in the network represent individual proteins, and the edges (lines connecting the nodes) indicate interactions between these proteins. The values along the edges denote the interaction confidence scores, which range from 0 to 1, reflecting the strength or likelihood of an interaction between each pair of proteins. From the network, we observe several high-confidence interactions. CLIP1 and CLASP2 exhibit an extremely high interaction score of 0.993, suggesting a robust functional relationship. This could indicate that these two proteins are part of a shared pathway or play complementary roles in cellular processes. TTN interacts with multiple proteins, such as KMT2C (interaction score 0.504), TPM1 (score 0.794), and SYNE2 (score 0.506). This suggests that TTN may have a central role in connecting multiple proteins, indicating a potential influence on several pathways in cancer biology. Another notable interaction is between CLIP1 and MYO1B, with an interaction score of 0.472, which suggests a moderate but biologically significant relationship. The PPI network provides insights into the complex biological interactions and the functional relationships between these proteins, potentially revealing collaborative roles in breast cancer pathogenesis. The intense interaction between CLIP1 and CLASP2 might indicate co-regulated functions in cytoskeletal dynamics or cellular structure maintenance, crucial in cancer metastasis and tumor growth. Similarly, the multiple connections of TTN suggest that it may act as a hub protein, influencing or coordinating multiple signaling pathways. The network helps visualize how individual proteins, deemed important by the XGBoost model, may work together or influence each other within the broader context of tumor biology. Understanding these connections can help uncover key mechanisms underlying tumor progression and may provide insights for targeted therapeutic interventions considering these relationships.
@@ 3.2 Clustering
Figure 3 illustrates a dendrogram resulting from the ensemble clustering of the patient dataset combined with hierarchical clustering for visualization. This dendrogram visually represents the grouping of patients, emphasizing the nested structure inherent in hierarchical clustering. After applying ensemble K-means clustering, a consensus (co-occurrence) matrix is constructed. The consensus matrix represents how often each pair of patients ends up in the same cluster across all ensemble runs. This essentially provides the stability and reliability of clustering by revealing consistent groupings. Once the consensus matrix is constructed, we use hierarchical clustering to create a more interpretable view of the relationships between different patients. The clustering is based on the similarity information contained in the consensus matrix. The dendrogram shows distinct and well-defined significant branches, indicating clear separations between different clusters within the dataset. Smaller sub-branches within the major clusters also point to additional layers of finer subgroupings, indicating subtle differences among patients. The hierarchical structure is apparent, with the merging heights reflecting similarity levels-patients with shorter linkage distances share more remarkable similarities.
We used co-occurrence matrices to visually represent the relationships between the three breast cancer markers across different clusters in Figure 4. In the first matrix, ER and PR status frequently occur together, indicating a moderate correlation between these markers. In contrast, ER status and HER2 status and PR status and HER2 status tend to appear in opposite states, suggesting a negative relationship between these markers. In the second one, ER and PR status are perfectly correlated, meaning they are always present in the same state (True or False) within this group. ER and HER2 status, along with PR and HER2 status, also tend to occur together, but this correlation is less consistent compared to ER and PR. Finally, ER and PR status exhibit a strong correlation in the third matrix, often appearing in the same state. However, ER and HER2 status and PR and HER2 status show minimal to no correlation, suggesting their occurrences are mainly independent of each other within this group.
Figure 3: Dendrogram resulting from the ensemble clustering of the patient dataset combined with hierarchical clustering
(a) Group 1 Co-Occurrence Matrix
(b) Group 2 Co-Occurrence Matrix
(c) Group 3 Co-Occurrence Matrix
Figure 4: Co-occurrence matrices for three different clusters of patients showing the correlations between the breast cancer markers.
@@ 4 Discussion
The relationships between ER/PR and HER2 are more complex, often showing a tendency for negative correlation. ER-positive and HER2-positive statuses are often mutually exclusive because they represent different molecular subtypes of breast cancer. ER/PR-positive tumors are usually classified as "luminal" subtypes, which tend to be more hormone-dependent and less aggressive compared to HER2-positive subtypes [25]. HER2-positive breast cancers tend to grow independently of hormone signaling. Instead, they rely on overexpression of the HER2 protein, which promotes cell growth via the MAPK and PI3K/AKT pathways. This reliance on a different growth mechanism often leads to a negative relationship between HER2 and hormone receptors [25]. In the first cluster, ER and HER2, as well as PR and HER2, often appear in opposite states. As shown in the Figure 4, the first cluster likely represents a more typical distribution of breast cancer subtypes, where hormone receptor-positive tumors (ER/PR+) are generally separate from HER2+ tumors. The moderate correlation suggests some overlap, including a mix of luminal A (strong ER/PR+) and luminal B subtypes (which can have lower PR levels and sometimes HER2 positivity). In the second cluster, the perfect correlation between ER and PR could indicate a cluster primarily composed of luminal A breast cancers, which typically show both strong ER and PR expression. The variable relationship with HER2 may reflect a small subset of luminal B cancers or suggest that HER2-positive tumors in this cluster co-express ER/PR, although not consistently. The last cluster suggests ER and PR are generally linked, but HER2 is mainly independent. It could represent a group with almost exclusively hormone receptor-positive, HER2-negative tumors. These tumors may be purely luminal A subtype, where hormone receptor positivity is dominant, and HER2 expression is largely absent.
To check the proteins represented in the correlation matrix in Figure 1 and their potential connections to the ER, PR, and HER2 status in breast cancer, it is essential to explore their involvement in cellular mechanisms relevant to breast cancer’s onset, progression, and differentiation. KMT2C is involved in chromatin remodeling and can influence gene expression through histone modifications, affecting ER and PR expression levels. Mutations in KMT2C may impact hormone receptor signaling and contribute to breast cancer subtypes like luminal A or B [26]. GCN1 regulates protein synthesis during stress. While its direct effect on receptor status is unclear, it may affect hormone receptor expression indirectly through stress response pathways relevant to breast cancer survival [27]. CLASP2 regulates microtubule dynamics, which is crucial for cell division and migration. HER2-positive cancers exhibit increased motility and may leverage proteins like CLASP2 for rapid increase and metastasis [28]. AQR is part of the RNA splicing process. It may impact ER and PR expression by regulating alternative splicing, potentially contributing to the heterogeneity observed in breast cancers [29]. MYO1B, a motor protein, facilitates cell movement and intracellular transport. Its role is critical in the increased invasive potential typical of HER2-positive or triple-negative cancers [30]. MAP2 stabilizes microtubules, potentially affecting cell response to hormone receptor or HER2 signaling. Microtubule dynamics are integral to both hormone-driven and HER2-driven pathways [31]. TTN maintains the cellular structure, which may be linked to ER or PR-positive luminal cancers with a more structured phenotype than HER2-enriched or triple-negative types [32]. CLIP1 helps regulate microtubule stability, vital for HER2-driven breast cancers that depend on rapid intracellular trafficking and signaling [33]. SYNE2 connects the cytoskeleton to the nucleus, affecting nuclear positioning and signaling and possibly influencing ER and PR effects. It could play a role in luminal subtypes with higher structural stability [34]. TPM1 stabilizes actin filaments, essential for cell shape and hormone receptor localization. It may help distinguish luminal and HER2-enriched cancers through cytoskeletal stability [35].
PPIs in breast cancer often indicate shared biological pathways that regulate critical cellular functions, including transcription regulation, cytoskeletal dynamics, and cell motility, affecting the expression and function of ER, PR, and HER2. Specifically, PPIs among proteins in the dataset point towards a collective role in coordinating receptor signaling and influencing breast cancer subtypes [36]. One interaction cluster involves KMT2C, TTN, SYNE2, and TPM1, all playing roles in chromatin structure, cellular integrity, and transcription regulation, impacting ER/PR expression and stabilization. KMT2C modifies chromatin to regulate gene expression, while TTN and SYNE2 contribute to structural integrity and nuclear positioning, essential for ER and PR-positive cancers [26, 32, 35]. TPM1 stabilizes actin filaments, supporting cellular shape and efficient receptor localization, further aiding in the regulation of ER and PR activity [35]. These proteins ensure the appropriate transcriptional environment and physical stability for hormone receptor functionality in luminal breast cancer subtypes. The cluster involving CLIP1, CLASP2, and MYO1B highlights their contribution to cytoskeletal dynamics and intracellular transport, which are crucial for receptor function and breast cancer cell motility. CLIP1 and CLASP2 stabilize microtubules and assist in receptor trafficking to the nucleus, supporting the activity of ER and PR [33, 28]. MYO1B is essential for actin-based motility and intracellular transport, supporting the increased invasive behavior characteristic of HER2-positive cancers [30]. These interactions facilitate efficient signaling, the key to both luminal and HER2-positive breast cancer subtypes. The PPIs suggest that different groups of proteins may serve as biomarkers or therapeutic targets for specific breast cancer subtypes. By understanding the functional modules within the PPI network, targeted therapeutic interventions can be developed for specific breast cancer subtypes, such as inhibiting CLASP2 and MYO1B to reduce invasion and metastasis in HER2-positive cancers.
@@
Integrating large language models like ProtGPT2 into bioinformatics allows researchers to capture nuanced protein sequence features, combining these with expression levels to create contextually relevant embeddings. In bioinformatics, protein expression levels are critical for determining a protein’s biological impact in a cell or tissue. Mathematically, this involves multiplying the embedding vector, E, of a protein sequence generated by ProtGPT2 by the corresponding protein expression level ($\lambda_{p_{n}}$), resulting in a new scaled embedding. This operation effectively scales each element of the embedding vector based on the protein’s expression level, introducing biological significance directly into the numerical representation. This scaling can be seen as a feature-specific weight that adjusts the impact of the embedding in downstream analyses, such as clustering or classification models, ensuring biologically essential proteins are given appropriate emphasis in machine learning models. Gating mechanisms are used in neural networks to modulate information processing, as seen in LSTMs or attention-based models like transformers [37, 20]. Multiplying the protein embedding vector by the expression level acts like a gating mechanism, where the expression level $\lambda_{p_{n}}$ serves as a gate that either amplifies or reduces the influence of the embedding based on the protein’s biological importance. When the expression level is high, more information passes through, akin to a gate being "opened" wider for influential inputs. In contrast, a lower expression level reduces the impact of that protein.

### v2

@@ Abstract
Breast cancer’s complexity and variability pose significant challenges in understanding its progression and guiding effective treatment. This study aims to integrate protein sequence data with expression levels to improve the molecular characterization of breast cancer subtypes and predict clinical outcomes. Using ProtGPT2, a language model designed for protein sequences, we generated embeddings that capture the functional and structural properties of proteins sequence. These embeddings were integrated with protein expression level to form enriched biological representations, which were analyzed using machine learning methods like ensemble K-means for clustering and XGBoost for classification. Our approach enabled successful clustering of patients into biologically distinct groups and accurately predicted clinical outcomes such as survival and biomarkers status, achieving high performance metrics, notably an F1 score of 0.88 for survival and 0.87 for biomarkers status prediction. Feature importance analysis identified KMT2C, CLASP2, and MYO1B as key proteins involved in hormone signaling, cytoskeletal remodeling, and therapy resistance in hormone receptor-positive and triple-negative breast cancer, with potential influence on breast cancer subtype behavior and progression. Furthermore, protein-protein interaction networks and correlation analyses revealed functional interdependencies among proteins that may influence breast cancer subtype behavior and progression. These findings suggest that integrating protein sequence and expression data provides valuable insights into tumor biology and has significant potential to enhance personalized treatment strategies in breast cancer care.
@@ 1 Introduction
Breast cancer is still a major global health issue, making up a significant number of cancer cases and deaths among women [1]. Despite advancements in early detection and treatment, the complexity and diversity of breast cancer present challenges in understanding its progression, predicting outcomes, and creating effective treatment plans. The classification of breast cancer into various subtypes based on molecular and clinical features is a crucial aspect of guiding treatment decisions and improving patient outcomes. Mainly, the presence or absence of specific biological markers, such as Estrogen Receptor (ER), Progesterone Receptor (PR), and Human Epidermal Growth Factor Receptor 2 (HER2), serves as a foundation for categorizing breast cancer into luminal, HER2-enriched, or basal-like subtypes [2, 3]. These markers are essential in determining how a tumor may respond to hormone therapy or other targeted treatments, making their accurate classification vital for personalized medicine [4].
@@
Recent advances in proteomics and genomics have significantly enhanced the molecular understanding of breast cancer. Proteomics, the large-scale study of proteins, adds a critical layer to cancer research since proteins are the functional entities in cells that carry out biological processes [5]. In breast cancer, shifts in protein expression patterns can provide valuable insights into tumor behavior, disease progression, and potential therapeutic targets. For instance, ER, PR, or HER2 changes can indicate how aggressive a tumor might be and what treatment strategies may be most effective. Understanding these markers not only aids in diagnosis but also plays a crucial role in therapeutic decision-making [4, 6].
@@
One of the key innovations in this work involves integrating protein expression data with the underlying protein sequences. Protein expression levels reflect the abundance of specific proteins within cells, shaped by processes such as translation and degradation, which are critical in regulating cellular functions. These dynamics play a significant role in determining cellular behavior, including in tumor contexts, where protein turnover and regulation influence functional outcomes and adaptations [7]. On the other hand, protein sequences, through their structural determinants, provide insights into functional domains, enabling a deeper understanding of their roles in biological processes. This sequence-structure-function relationship forms the foundation for exploring how specific protein functions emerge, offering a pathway to study processes such as cancer progression by integrating structural and functional annotations [8]. Combining these two layers of information, protein expression level, and sequence features, we can capture both quantitative and qualitative aspects of the proteins, leading to a more comprehensive understanding of their roles in breast cancer biology.
@@
We used ProtGPT2 [9], an advanced large language model (LLM) based on the GPT-2 architecture, to facilitate the integration of protein sequence and expression data. ProtGPT2 is trained explicitly for protein sequences, enabling it to generate meaningful embeddings that represent proteins’ functional and structural properties. These embeddings provide a high-dimensional representation that captures complex details of each protein sequence, allowing the model to leverage both local and global features of the sequences [9]. Integrating transcriptomic data with protein expression levels in breast cancer enables the construction of a biologically comprehensive representation of tumor biology, facilitating downstream analyses such as clustering tumors into proteomic subgroups and associating these clusters with patient survival outcomes and molecular subtypes [10]. Our approach, which incorporates machine learning methods like ensemble K-means for clustering and classification algorithms such as XGBoost and Random Forest, aims to uncover relationships between protein features and clinical characteristics, offering new insights into the grouping of breast cancer patients and potential predictive biomarkers. Furthermore, the importance of trust and explainability in using artificial intelligence (AI) for healthcare has been highlighted in recent studies, including [11] that explore mechanisms to bridge the gap between advanced AI technologies and clinical practice, specifically in the breast cancer. In this context, the integration of protein sequence embeddings with expression levels may be considered a biologically meaningful framework for understanding protein function. Coupled with machine learning methods, enables the categorization of breast cancer patients and identification of potential predictive biomarkers.
@@ 2.1 Data Source
In our work, we utilized the dataset of 12,553 proteins from the TCGA proteomics and then extracted protein sequence data from the National Center for Biotechnology Information (NCBI) database [14]. The protein sequence data were retrieved using API calls based on the RefSeq IDs provided in the TCGA dataset. The extracted protein sequences were obtained in Fast-All (FASTA) format, ensuring compatibility with further bioinformatic and proteomic analyses. This additional data collection allowed for integrating protein sequence information with the proteomic expression levels, enriching the dataset used in our study.
@@ 2.5 Classification
Different classification algorithms such as XGBoost and Random Forest are applied to classify the patient data to predict different aspects like survival of patients. One of these predictions is to predict the status of three biological markers, ER, PR status, and HER2 Final Status, which are essential in breast cancer classification. Each marker is a target variable, a multi-output classification problem. The $X_{m}$ is used to predict them as a multi-label. The Hamming loss metric measures the fraction of incorrectly predicted labels, providing an overall view of the model’s ability to classify each target accurately. Lower Hamming loss indicates better classification performance. The formula for Hamming loss is given by [18]:
@@ 3.1 Classifications
To evaluate the effectiveness and robustness of integrating protein sequence data with protein expression levels, we applied different machine learning algorithms to predict the survival outcomes of breast cancer patients. The survival status, defined as either "alive" or "deceased", was predicted based on a comprehensive feature set comprising the integrated protein sequence and expression level and relevant clinical data such as tumor size and patient age. The classification model demonstrated a strong performance on the XGBoost, achieving an F1 score of 0.88. This result underscores the model’s ability to distinguish between patients who are likely to survive and those at a higher risk.
Then, We employed multiple classification algorithms to evaluate the predictive capability of our integrated protein sequence and expression level in determining the status of breast cancer biomarkers, including ER, PR, and HER2 status. We treated the prediction of these three markers as a multi-output classification problem, where each marker represents an individual target to be predicted simultaneously. Table 1 compares various classification algorithms, showcasing the precision, recall, F1 score, and Hamming loss for each model. These metrics are reported as weighted averages, reflecting the model’s ability to predict multiple target labels across all patient samples effectively.
@@
Table 1: Precision, recall, and F1 score for predicting all markers as a multi-output classification problem based on different algorithms reported in the weighted average
The results in Table 1 show that XGBoost achieved the best overall performance, with an F1 score of 0.87 and balanced precision and recall values of 0.88 each. This indicates that XGBoost was highly influential in predicting the multi-output targets accurately, achieving strong generalization across all three biomarkers while maintaining a relatively low Hamming loss of 0.18. CatBoost and AdaBoost also performed well, with CatBoost showing a slight edge in precision and Hamming loss compared to XGBoost.
To further investigate which proteins were most influential in predicting the biomarker statuses, we analyzed the feature importance scores from the XGBoost model. The integration of protein expression levels and sequence information was used for each protein, allowing the model to leverage both the abundance of the proteins and their functional sequence-derived attributes. Table 2 lists the top 10 proteins, including critical proteins with the highest importance scores during classification. Among the proteins, KMT2C and GCN1 were the most influential proteins, with importance scores of 0.119 and 0.114, respectively.
@@
Table 2: Importance score values for top 10 important proteins from XGBoost model
@@
Figure 1: Pearson correlation matrix of the top 10 important proteins in the biomarkers prediction
Figure 2 presents the protein-protein interaction (PPI) network for the top 10 most important proteins identified from the XGBoost feature importance analysis. This network was constructed using STRING-db [19], a well-known database for retrieving known and predicted protein-protein interactions. The nodes in the network represent individual proteins, and the edges (lines connecting the nodes) indicate interactions between these proteins. The values along the edges denote the interaction confidence scores, which range from 0 to 1, reflecting the strength or likelihood of an interaction between each pair of proteins. From the network, we observe several high-confidence interactions. CLIP1 and CLASP2 exhibit an extremely high interaction score of 0.993, suggesting a robust functional relationship. This could indicate that these two proteins are part of a shared pathway or play complementary roles in cellular processes. TTN interacts with multiple proteins, such as KMT2C (interaction score 0.504), TPM1 (score 0.794), and SYNE2 (score 0.506). This suggests that TTN may have a central role in connecting multiple proteins, indicating a potential influence on several pathways in cancer biology. Another notable interaction is between CLIP1 and MYO1B, with an interaction score of 0.472, which suggests a moderate but biologically significant relationship. The PPI network provides insights into the complex biological interactions and the functional relationships between these proteins, potentially revealing collaborative roles in breast cancer pathogenesis.
@@ 3.2 Clustering
Figure 3 presents the dendrogram resulting from ensemble clustering of the patient dataset, with hierarchical clustering applied for visualization. The dendrogram reveals a clear hierarchical structure, separating patients into three major color-coded clusters: orange, green, and red. These clusters indicate well-defined patient subgroups based on consensus similarity patterns, with nested sub-branches reflecting additional fine-scale subgroupings. Shorter linkage distances between branches represent higher similarity, while the distinct separation between color groups underscores strong intergroup differences.
To further explore the biological relevance of these clusters, we analyzed the co-occurrence of three key breast cancer markers of ER, PR, and HER2 within each group, as shown in Appendix Figure A.1. In the orange cluster, ER and PR status frequently occur together, indicating a moderate correlation between these markers. In contrast, ER status and HER2 status and PR status and HER2 status tend to appear in opposite states, suggesting a negative relationship between these markers. In the green cluster, ER and PR status are perfectly correlated, meaning they are always present in the same state (True or False) within this group. ER and HER2 status, along with PR and HER2 status, also tend to occur together, but this correlation is less consistent compared to ER and PR. Finally, the red cluster shows ER and PR status exhibit a strong correlation, often appearing in the same state. However, ER and HER2 status and PR and HER2 status show minimal to no correlation, suggesting their occurrences are mainly independent of each other within this group.
Figure 3: Dendrogram resulting from the ensemble clustering of the patient dataset combined with hierarchical clustering for each patient case
@@ 4 Discussion
The relationships between ER/PR and HER2 are more complex, often showing a tendency for negative correlation. ER-positive and HER2-positive statuses are often mutually exclusive because they represent different molecular subtypes of breast cancer. ER/PR-positive tumors are usually tend to be more hormone-dependent and less aggressive compared to HER2-positive subtypes [21]. HER2-positive breast cancers tend to grow independently of hormone signaling. Instead, they rely on overexpression of the HER2 protein, which promotes cell growth via the MAPK and PI3K/AKT pathways [22]. This reliance on a different growth mechanism often leads to a negative relationship between HER2 and hormone receptors [21]. In the first cluster, ER and HER2, as well as PR and HER2, often appear in opposite states. As shown in Appendix Figure A.1, In the first cluster, ER and PR are positively correlated, while both show moderate negative correlations with HER2 expression. This inverse relationship between hormone receptors and HER2 is consistent with immunohistochemistry (IHC) based breast cancer subtypes, where ER/PR+ tumors (Luminal A and B) are generally distinct from HER2-enriched or triple-negative tumors [21]. In the second cluster, the perfect correlation between ER and PR could indicate a cluster primarily composed of luminal A breast cancers, which typically show both strong ER and PR expression. The variable relationship with HER2 may reflect a small subset of luminal B cancers or suggest that HER2-positive tumors in this cluster co-express ER/PR, although not consistently. The last cluster suggests ER and PR are generally linked, but HER2 is mainly independent. It could represent a group with almost exclusively hormone receptor-positive, HER2-negative tumors. These tumors may be purely luminal A subtype, where hormone receptor positivity is dominant, and HER2 expression is largely absent. The analysis reveals an overall inverse relationship between hormone receptors (ER/PR) and HER2, consistent with known breast cancer subtypes like Luminal A and HER2-enriched. Notably, the second cluster shows perfect ER/PR correlation with variable HER2 expression, which may represent a hybrid or unusual luminal B subgroup. This pattern suggests a potential novel cluster characterized by co-expression of hormone receptors and moderate HER2, warranting further investigation into its molecular features and clinical behavior.
[…]
PPIs in breast cancer often indicate shared biological pathways that regulate critical cellular functions, including transcription regulation, cytoskeletal dynamics, and cell motility, affecting the expression and function of ER, PR, and HER2. Specifically, PPIs among proteins in the dataset point towards a collective role in coordinating receptor signaling and influencing breast cancer subtypes [31]. The interaction network reveals two distinct functional modules. The first, centered on TTN, connects chromatin remodeling (via KMT2C) with cytoskeletal stability (TPM1 and SYNE2), suggesting a nuclear structural axis that may regulate hormone receptor function and support luminal subtype differentiation [23, 30, 29]. The second module, linking CLIP1, CLASP2, and MYO1B, highlights a microtubule motor axis associated with migration, EMT, and potential therapy resistance [25, 27], may particularly in HER2-positive and triple-negative breast cancers. The strong interaction between CLIP1 and CLASP2 (interaction score: 0.993) suggests tightly coordinated roles in cytoskeletal remodeling and structural maintenance, which are critical for tumor progression and metastasis [25]. This network highlights how individual proteins, prioritized by the XGBoost model, may functionally converge or influence each other within the broader landscape of tumor biology. Understanding these interactions provides valuable insights into the mechanisms driving tumor progression and may inform the development of targeted therapeutic strategies that account for these molecular relationships.
@@
Integrating large language models like ProtGPT2 into bioinformatics allows researchers to capture nuanced protein sequence features, combining these with expression levels to create contextually relevant embeddings. In bioinformatics, protein expression levels are critical for determining a protein’s biological impact in a cell or tissue. Mathematically, this involves multiplying the embedding vector, E, of a protein sequence generated by ProtGPT2 by the corresponding protein expression level ($\lambda_{p_{n}}$), resulting in a new scaled embedding. This operation effectively scales each element of the embedding vector based on the protein’s expression level, introducing biological significance directly into the numerical representation. This scaling can be seen as a feature-specific weight that adjusts the impact of the embedding in downstream analyses, such as clustering or classification models, ensuring biologically essential proteins are given appropriate emphasis in machine learning models. Gating mechanisms are used in neural networks to modulate information processing, as seen in LSTMs or attention-based models like transformers [32, 16]. Multiplying the protein embedding vector by the expression level acts like a gating mechanism, where the expression level $\lambda_{p_{n}}$ serves as a gate that either amplifies or reduces the influence of the embedding based on the protein’s biological importance. When the expression level is high, more information passes through, akin to a gate being "opened" wider for influential inputs. In contrast, a lower expression level reduces the impact of that protein. Using this biologically-informed embedding strategy, our method produced promising results in both clustering and classification tasks. Notably, it enabled the identification of biologically relevant protein clusters and revealed potential subtype-specific protein signatures, demonstrating its utility in uncovering meaningful patterns within complex omics data.